Find the change in volume of 30 L of a fluid when subjected to a


pressure increase of 2.0 MPa. The bulk modulus of elasticity of the


fluid is 2.24 GPa.

1
Expert's answer
2022-01-20T13:50:46-0500

Answer

pressure increase of P=2.0 MPa.

Volume V=30L

The bulk modulus of elasticity of the

fluid is K=2.24 GPa.

So change in volume is given by

"\Delta V=\frac{PV}{K}"

Putting all values

"\Delta V=\frac{2.0\times10^6\times30\times10^{-3}}{2.24\times10^9}\\=2.68\times10^{-5}\\=0.0268L"